OpenStudy (anonymous):

The table below shows the cube roots of different numbers Number(x) 8 27 64 125 Cube root(y) 2 3 4 5 Part A: Does the table represent y as a function of x? Justify your answer. (5 points) Part B: The total cost f(x), in dollars, for renting a bike for x hours is shown below: f(x) = 10 + 20x What is the value of f(100), and what does f(100) represent? (5 points)

OpenStudy (anonymous):

for b I did 10 + 20x 10+ 20(100) 10+2000 2010

OpenStudy (anonymous):

A. Yes. y is a function of x if there are not two different y values for the same x value. B. 10+ 20times100=2000+10=2010 f(100) represents an x value to plug into the equation.

OpenStudy (anonymous):

good job!

OpenStudy (anonymous):

thanks for the help so quickly

OpenStudy (anonymous):

no problem. i just happened to see you question pop up after my question was answered, so I thought why not? Always happy to help a brudda out! (if you're a brudda. If you're not, it was good to help a sista out!)